# NILE [Uncaptioned image] : Internal Consistency Alignment in Large Language Models

### v1

@@ A.1.3 IKE Sample Demonstration
Instruction $\mathrm{instruction}^{o}$: Recommend 3 books that could get someone into a new field of study. Input $\mathrm{input}^{o}$: None 2-shot Demonstration $f^{demo}(q^{o})=\{(q^{\mathcal{R}}_{1},ik^{\mathcal{R}}_{1}),(q^{\mathcal{R}}_{2},ik^{\mathcal{R}}_{2})\}$: $\mathbf{q^{\mathcal{R}}_{1}}:$ Provide 3 pieces of advice for someone who is struggling to stay motivated in their studies. $\mathbf{ik^{\mathcal{R}}_{1}}:$ 1. Understanding Motivation: It is essential to comprehend the concept of motivation, including intrinsic and extrinsic motivation. Intrinsic motivation refers to doing something because it is inherently interesting or enjoyable, while extrinsic motivation involves doing something because it leads to a separable outcome. Understanding these types of motivation can help identify what drives an individual’s desire to study and how to bolster it. 2. Goal Setting: Knowledge of effective goal-setting strategies is crucial. Setting specific, measurable, achievable, relevant, and time-bound (SMART) goals can help create a clear roadmap for studying. This process can help break down the overwhelming feeling into manageable tasks, making it easier to stay motivated. …… $\mathbf{q^{\mathcal{R}}_{2}}:$ Please recommend a movie for someone who likes animated films. $\mathbf{ik^{\mathcal{R}}_{2}}:$ 1. Understanding Animated Films: Animated films, also known as cartoons, use drawn or computer-generated imagery to create the illusion of movement. They can be in 2D, 3D, stop-motion or other animation techniques. Knowledge of the different styles and techniques of animation is crucial to recommending films that align with specific preferences. 2. Popular Animation Studios: Some of the most prominent animation studios include Disney, Pixar, DreamWorks, Studio Ghibli, and Illumination. Each studio has its own unique style and storytelling approach. Familiarity with the filmographies of these studios can help in making informed recommendations. 3. Genres within Animation: Animated films encompass a wide range of genres, from adventure and fantasy to drama and comedy. Some animated films are targeted towards children, while others might appeal to a broader age range including adults. Understanding the various genres and their target audiences can guide the recommendation process. 4. Cultural and Thematic Elements: Animated films often incorporate diverse cultural stories and themes. Films might explore complex themes like identity, family, and morality, or might be more light-hearted and comedic. Awareness of these elements can help match a film to a viewer’s personal interests and values. ……
@@ A.1.4 Knowledge-aware Sample Revision (KSR)
For KSR, we also use gpt-4-turbo-2024-04-09 endpoint as revisor agent $\mathcal{A}_{r}$. We use openai 1.42.0 python library with n set to 1, temperature to 0.7, and max_tokens to 1,024. We run KSR on 52,000 samples from the Alpaca dataset and 50,000 samples from the OpenOrca dataset. Case studies of KSR can be found in Table 13. These results display the KSR’s capability to infuse internal knowledge information into original answers through revision.
@@ A.1.5 Sample Revision (SR)
Unlike KSR, Sample Revision (SR) does not revise for each instruction pair $(q^{o},a^{o})$. Therefore, $\mathcal{A}_{r}$ in SR only uses external knowledge such as world knowledge from $(q^{o},a^{o})$ and its own parameter knowledge, being completely isolated from internal knowledge $ik^{\mathcal{M}}_{q^{o}}$ of $\mathcal{M}$. Table 11 shows the detailed prompt of the revisor $\mathcal{A}_{r}$ in SR.
Provide a better response based on "$\mathbf{\{a^{o}\}}$" to comply with given instruction, input, and related knowledge. Instruction: $\mathbf{\{\mathrm{\textbf{instruction}}^{o}\}}$ Input:$\mathbf{\{\mathrm{\textbf{input}}^{o}\}}$ Please directly output the improved response.
@@
A.2 Experiment Details
A.2.1 Benchmarks
We use the officially recommended settings from all benchmarks for evaluation. For Alpaca-Eval V2, we use "alpaca_eval_cot_gpt4_turbo_fn" as annotators, and we set max_new_tokens to 1024, temperature to 1.0, top_p to 1.0 and batch_size to 128. For Arena-Hard, we set temperature to 0.0, max_tokens to 1024, judge_model to gpt-4-1106-preview, baseline_model to gpt-4-0314, and num_choices to 1. For BBH and MTBench, we use the default settings in the official source code.
@@ A.2.1 Benchmarks
A.2.2 Fixed Demonstration (FD)
Instruction: "Give three tips for staying healthy." Related Knowledge: Importance of health: Maintaining good health is crucial for overall well-being and quality of life. * Factors affecting health: A person’s health can be influenced by various factors such as diet, exercise, sleep, stress, and genetics. * Prevention is key: Preventing illnesses and maintaining good health requires adopting healthy habits and making lifestyle changes. * Healthy habits: Developing healthy habits such as regular exercise, balanced diet, and adequate sleep can help prevent chronic diseases. * Lifestyle changes: Making lifestyle changes such as quitting smoking, reducing sugar intake, and managing stress can also contribute to good health. * Importance of self-care: Taking care of one’s physical, emotional, and mental health is essential for overall well-being. * Access to healthcare: Having access to quality healthcare and medical facilities is also important for staying healthy. * Healthy behaviors: Engaging in healthy behaviors such as regular check-ups, vaccinations, and screenings can help prevent illnesses and detect health issues early. * Health literacy: Having knowledge and understanding of health-related information is important for making informed decisions about one’s health. Instruction: "What are the three primary colors?" Related Knowledge: Primary colors are colors that cannot be created by mixing other colors together. * The three primary colors are: 1. Red 2. Blue 3. Yellow * Primary colors are the base colors used to create all other colors. * By mixing different combinations of primary colors, you can create secondary colors, tertiary colors, and a wide range of shades and hues. * Primary colors are often used in art, design, and painting to create bold and vibrant colors. * The primary colors are also used in color theory to understand how colors interact with each other and how they can be used to create contrast, harmony, and balance. * In addition to art and design, primary colors are also used in science, technology, engineering, and mathematics (STEM) fields, such as physics, chemistry, and biology, to describe and analyze the properties of light and color. * The primary colors are a fundamental concept in many cultures and have been used in art and design for centuries, with examples found in ancient civilizations such as Egypt, Greece, and China. Instruction: $\mathbf{\{\mathrm{\textbf{instruction}}^{o},\mathrm{\textbf{input}}^{o}\}}$ Related Knowledge:
@@ A.2.2 Fixed Demonstration (FD)
A.2.3 Effects of KSR
Sentence embedding similarity distribution plots across datasets and models are Table 4. The average similarity score of KSR outputs in these plots consistently exceeds that of the other two outputs. All distribution are proven by the Chi-squared Test to be significantly different from each other with the p-value lower than 0.01.
@@ A.2.3 Effects of KSR
Vanilla: "Reminder: Renew Your Subscription Today!"
Revised: "Action Required: Time to Renew Your Subscription!"
@@
Hyperbole is a figure of speech that uses exaggeration for emphasis or effect.* Hyperbolic metaphors are metaphors that use hyperbole to create a vivid and dramatic image.* For example, "The crowd was so large that it stretched from one end of the stadium to the other."* This metaphor uses hyperbole to create a dramatic image of the crowdś size.* Hyperbolic metaphors can be used to create a strong emotional response or to make a point more effectively.* They can also be used to add humor or to create a sense of exaggeration or absurdity.
Create a 3-5 sentence definition for the term "data science".

### v2

@@ A.1.3 IKE Sample Demonstration
Instruction $\mathrm{instruction}^{o}$: Recommend 3 books that could get someone into a new field of study.
Input $\mathrm{input}^{o}$: None
2-shot Demonstration $f^{demo}(q^{o})=\{(q^{\mathcal{R}}_{1},ik^{\mathcal{R}}_{1}),(q^{\mathcal{R}}_{2},ik^{\mathcal{R}}_{2})\}$:
$\mathbf{q^{\mathcal{R}}_{1}}:$ Provide 3 pieces of advice for someone who is struggling to stay motivated in their studies. $\mathbf{ik^{\mathcal{R}}_{1}}:$ 1. Understanding Motivation: It is essential to comprehend the concept of motivation, including intrinsic and extrinsic motivation. Intrinsic motivation refers to doing something because it is inherently interesting or enjoyable, while extrinsic motivation involves doing something because it leads to a separable outcome. Understanding these types of motivation can help identify what drives an individual’s desire to study and how to bolster it. 2. Goal Setting: Knowledge of effective goal-setting strategies is crucial. Setting specific, measurable, achievable, relevant, and time-bound (SMART) goals can help create a clear roadmap for studying. This process can help break down the overwhelming feeling into manageable tasks, making it easier to stay motivated. …… $\mathbf{q^{\mathcal{R}}_{2}}:$ Please recommend a movie for someone who likes animated films. $\mathbf{ik^{\mathcal{R}}_{2}}:$ 1. Understanding Animated Films: Animated films, also known as cartoons, use drawn or computer-generated imagery to create the illusion of movement. They can be in 2D, 3D, stop-motion or other animation techniques. Knowledge of the different styles and techniques of animation is crucial to recommending films that align with specific preferences. 2. Popular Animation Studios: Some of the most prominent animation studios include Disney, Pixar, DreamWorks, Studio Ghibli, and Illumination. Each studio has its own unique style and storytelling approach. Familiarity with the filmographies of these studios can help in making informed recommendations. 3. Genres within Animation: Animated films encompass a wide range of genres, from adventure and fantasy to drama and comedy. Some animated films are targeted towards children, while others might appeal to a broader age range including adults. Understanding the various genres and their target audiences can guide the recommendation process. 4. Cultural and Thematic Elements: Animated films often incorporate diverse cultural stories and themes. Films might explore complex themes like identity, family, and morality, or might be more light-hearted and comedic. Awareness of these elements can help match a film to a viewer’s personal interests and values. ……
@@ A.1.4 Knowledge-aware Sample Revision (KSR)
For KSR, we also use gpt-4-turbo-2024-04-09 endpoint as revisor agent $\mathcal{A}_{r}$. We use the OpenAI 1.42.0 Python library with n set to 1, temperature to 0.7, and max_tokens to 1,024. We run KSR on 52,000 samples from the Alpaca dataset and 50,000 samples from the OpenOrca dataset. Case studies of KSR can be found in Table 19. These results display the KSR’s capability to infuse internal knowledge information into original answers through revision.
@@ A.1.5 Sample Revision (SR)
Unlike KSR, Sample Revision (SR) does not revise for each instruction pair $(q^{o},a^{o})$. Therefore, $\mathcal{A}_{r}$ in SR only uses external knowledge such as world knowledge from $(q^{o},a^{o})$ and its own parameter knowledge, being completely isolated from internal knowledge $ik$ of $\mathcal{M}$. Table 12 shows the detailed prompt of the revisor $\mathcal{A}_{r}$ in SR.
[…]
@@
A.1.6 Statistics of Inference Overhead
In the IKE step of NILE, internal knowledge is efficiently extracted from sample datasets within 6 hours using vLLM on an 8-A6000 GPU server. To further show how much inference overhead is introduced, we measured the average token usage per sample for this step, which is detailed in Table 13.
Model / Dataset Llama Mistral
OpenOrca 684.8 801.1
Alpaca 620.6 544.5
Table 13: Average generated tokens per sample using vLLM during IKE across different datasets and models.
The Knowledge-aware Sample Revision (KSR) step further optimizes efficiency, with GPT-4 achieving modest token usage per sample, also shown in Table 14. This results in exceptionally low operational costs, making our approach scalable, cost-effective, and practical for real-world applications.
Model / Dataset Llama Mistral
OpenOrca \$0.015/182.0 \$0.016/182.1
Alpaca \$0.013/193.7 \$0.011/165.3
Table 14: Price (in USD) and average generated tokens per sample during KSR across different datasets and models.
A.2 Evaluating NILE’s Effectiveness on External Knowledge-Intensive Tasks
To briefly examine some of NILE’s potential issues and limitations, we conducted an additional experiment on the SQuADv2 Rajpurkar et al. (2018) validation set using our sampled Alpaca-GPT4 dataset under the same settings outlined in our paper. The SQuADv2 validation set was chosen because it contains 119,000 test samples of reading comprehension, where large language models (LLMs) must answer questions based on external knowledge provided in corresponding supporting passages. As such, it serves as a suitable and rigorous benchmark for evaluating an LLM’s ability to comprehend and utilize external knowledge effectively. The results in Table 16 demonstrate that NILE can positively influence this capability.
A.3 Examining NILE’s Performance on Multitask Accuracy
In order to further test the extensiveness of NILE’s improvement on LLMs, we have conducted additional experiments on MMLU using two models (Meta-LLama-3.1-8B and Mistral-7B-v0.3) and two IFT datasets (Alpaca and Orca). All experiments adhered to the official default configuration from the lm-evaluation-harness implementation [^18] https://github.com/EleutherAI/lm-evaluation-harness/tree/main/lm_eval/tasks/mmlu. The results presented in Table 15 demonstrate that NILE consistently achieves significant performance improvements in highly complex QA settings like MMLU. Interestingly, we observed a noticeable dip in accuracy with the SR baseline across both datasets and models. This result further underscores the necessity of incorporating internal knowledge within the NILE framework to enhance alignment in IFT datasets.
Method Accuracy (%) $\uparrow$
Mistral-7b-v0.3
Alpaca Vanilla 57.21
Alpaca + SR 56.25
Alpaca + NILE 57.56
Orca Vanilla 56.92
Orca + SR 54.54
Orca + NILE 56.91
Meta-Llama-3.1-8B
Alpaca Vanilla 62.51
Alpaca + SR 62.41
Alpaca + NILE 63.93
Orca Vanilla 62.68
Orca + SR 62.19
Orca + NILE 63.09
Table 15: Experiment results of NILE on MMLU benchmark. The highest values are bolded.
Method EM $\uparrow$ F1 $\uparrow$
Mistral-7b-v0.3
Alpaca vanilla 4.91 14.13
Alpaca + NILE 5.61 14.51
Meta-Llama-3.1-8B
Alpaca vanilla 1.41 9.35
Alpaca + NILE 4.51 11.65
Table 16: Experiment results of NILE on SQuADv2 dataset. The highest values are bolded.
Meta-Llama-3.2-3B
Alpaca vanilla 3.50 6.17 / 3.54 5.52 43.63
Alpaca + SR 3.00 6.18 / 4.43 5.65 43.02
Alpaca + NILE 4.60 6.61 / 5.10 5.94 43.74
Orca vanilla 3.60 5.61 / 4.41 2.20 43.36
Orca + SR 4.00 6.18 / 5.51 3.06 41.87
Orca + NILE 4.20 10.77 / 8.46 4.63 45.69
Qwen2.5-7B
Alpaca vanilla 11.90 13.84 / 7.83 7.13 45.95
Alpaca + SR 14.40 15.40 / 8.90 6.78 45.94
Alpaca + NILE 18.40 17.42 / 12.24 8.13 45.72
Orca vanilla 14.90 19.19 / 13.42 6.60 46.89
Orca + SR 13.60 18.45 / 14.34 7.05 49.32
Orca + NILE 17.10 20.55 / 16.58 7.31 48.61
Qwen2.5-14B
Alpaca vanilla 12.80 15.37 / 8.16 7.33 48.01
Alpaca + SR 20.00 21.56 / 12.59 7.73 48.69
Alpaca + NILE 22.60 28.55 / 17.45 8.06 49.01
Orca vanilla 24.00 19.99 / 17.44 7.68 48.49
Orca + SR 26.60 24.40 / 18.85 7.90 47.66
Orca + NILE 31.90 32.12 / 29.82 8.21 49.96
Table 17: Complete experiment results of more LLMs on Alpaca and OpenOrca datasets for Arena-Hard, Alpaca-Eval V2 LCWR, MTBench, and BBH benchmarks. The highest values are bolded, and the second highest is underlined.
A.4 Experiment Details
A.4.1 Benchmarks
We use the officially recommended settings from all benchmarks for evaluation. For Alpaca-Eval V2, we use ”alpaca_eval_cot_gpt4_turbo_fn” as annotators, and we set max_new_tokens to 1024, temperature to 1.0, top_p to 1.0, and batch_size to 128. For Arena-Hard, we set the temperature to 0.0, max_tokens to 1024, judge_model to gpt-4-1106-preview, baseline_model to gpt-4-0314, and num_choices to 1. For the Arena-Hard and Alpaca-Eval V2 benchmark, we keep the same alpaca-style system prompt as the fine-tuning stage during evaluation. As for BBH and MTBench, we use the default settings in the official source code.
@@ A.4.1 Benchmarks
A.4.2 Fixed Demonstration (FD)
Instruction: ”Give three tips for staying healthy.” Related Knowledge: * Importance of health: Maintaining good health is crucial for overall well-being and quality of life. * Factors affecting health: A person’s health can be influenced by various factors such as diet, exercise, sleep, stress, and genetics. * Prevention is key: Preventing illnesses and maintaining good health requires adopting healthy habits and making lifestyle changes. * Healthy habits: Developing healthy habits such as regular exercise, balanced diet, and adequate sleep can help prevent chronic diseases. * Lifestyle changes: Making lifestyle changes such as quitting smoking, reducing sugar intake, and managing stress can also contribute to good health. * Importance of self-care: Taking care of one’s physical, emotional, and mental health is essential for overall well-being. * Access to healthcare: Having access to quality healthcare and medical facilities is also important for staying healthy. * Healthy behaviors: Engaging in healthy behaviors such as regular check-ups, vaccinations, and screenings can help prevent illnesses and detect health issues early. * Health literacy: Having knowledge and understanding of health-related information is important for making informed decisions about one’s health. Instruction: ”What are the three primary colors?” Related Knowledge: * Primary colors are colors that cannot be created by mixing other colors together. * The three primary colors are: 1. Red 2. Blue 3. Yellow * Primary colors are the base colors used to create all other colors. * By mixing different combinations of primary colors, you can create secondary colors, tertiary colors, and a wide range of shades and hues. * Primary colors are often used in art, design, and painting to create bold and vibrant colors. * The primary colors are also used in color theory to understand how colors interact with each other and how they can be used to create contrast, harmony, and balance. * In addition to art and design, primary colors are also used in science, technology, engineering, and mathematics (STEM) fields, such as physics, chemistry, and biology, to describe and analyze the properties of light and color. * The primary colors are a fundamental concept in many cultures and have been used in art and design for centuries, with examples found in ancient civilizations such as Egypt, Greece, and China. Instruction: $\mathbf{\{\mathrm{\textbf{instruction}}^{o},\mathrm{\textbf{input}}^{o}\}}$ Related Knowledge:
@@ A.4.2 Fixed Demonstration (FD)
A.4.3 Effects of KSR
To validate KSR’s effectiveness in enhancing internal consistency between world knowledge from instructions and the model’s internal knowledge, we conducted experiments measuring the similarity between extracted internal knowledge and baseline knowledge across different models and datasets. In Llama-3 and Mistral, we used the instructions from the Alpaca and Orca as prompts to evaluate the models’ internal knowledge. Then, we obtained the models’ vanilla output for these instructions, the output adjusted using KSR, and the output using SR. We randomly sampled 10,000 instructions to calculate the sentence similarity between these three outputs and the internal knowledge. As demonstrated in Figure 4, the similarity score distribution of the outputs adjusted by KSR is significantly closer to 1 compared to the other two methods, with Chi-squared test p-values lower than 0.01.
These results provide strong evidence supporting the effectiveness of the KSR approach in increasing the internal consistency from instructions by integrating relevant world and internal knowledge. The case study of KSR is listed in Table 19.
@@ A.4.3 Effects of KSR
Vanilla: ”Reminder: Renew Your Subscription Today!”
Revised:Action Required: Time to Renew Your Subscription!”
@@
Hyperbole is a figure of speech that uses exaggeration for emphasis or effect.* Hyperbolic metaphors are metaphors that use hyperbole to create a vivid and dramatic image.* For example, ”The crowd was so large that it stretched from one end of the stadium to the other.”* This metaphor uses hyperbole to create a dramatic image of the crowdś size.* Hyperbolic metaphors can be used to create a strong emotional response or to make a point more effectively.* They can also be used to add humor or to create a sense of exaggeration or absurdity.
Create a 3-5 sentence definition for the term ”data science”.